How many grams of silver chloride are formed when 11.0 g of silver nitrate reacts with 14.0 g of barium chloride
1
Expert's answer
2020-06-19T07:01:52-0400

"2AgNO_3(aq)+BaCl_2(aq)=Ba(NO_3)_2(aq)+2AgCl(s)"

moles of "AgNO_3 = 11.0 g \frac{1mole}{169.87g}=0.0648 mol"

Moles of "BaCl_2= 14.0 g \frac{1 mole}{208.23 g}= 0.0672 mol"

Find limiting reactant:

Compare "\frac{moles AgNO_3}{2}" and "\frac{moles BaCl_2}{1}"

"\frac{0.0648}{2}<\frac{0.0672}{1}" AgNO3 is limiting reactant.

Find mass of AgCl

"0.0648 mol AgNO_3 \frac{ 2 mol AgCl}{2 mol AgNO_3} \frac{143.32 g AgCl}{1 mol AgCl}= 9.29 g AgCl"